# Quantitative surgery and total mean curvature

### v1

@@ 7. A new quasi-local mass
Theorem 1.2 has important applications in Physics. In 1982, Penrose [23] listed important open problems in General Relativity. Finding a suitable notion of quasi-local mass was listed first.
@@
In order to define the mass of a finite region–known as quasi-local mass–various proposals have been made, including the Hawking mass [19], the Bartnik mass [6], the Brown–York mass [13], and the Wang–Yau mass [28]. Each of these notions has attractive geometric features but also important limitations. A striking common feature is that all of them are particularly well-suited for spherical topology: The Hawking mass often underestimates the true mass and is most effectively used together with inverse mean curvature flow, whose monotonicity holds only for $M\cong S^{2}$. The Brown–York and Wang–Yau masses rely on isometric embeddings into Euclidean or Minkowski space, and classical embedding results such as the Weyl theorem require positive Gauss curvature. The Bartnik mass, perhaps the most mathematically natural proposal, has so far been explicitly computed only for spherical boundaries.
This contrasts with the increasing importance of spacetimes with more complicated topology including Emparan–Reall’s black ring [16], the Kottler metric [21], and the Horowitz–Myers soliton [11].
Motivated by [22] we define a new notion of quasi-local mass which has no such topological shortcomings. For this purpose, let $(N,g_{N})$ be a complete manifold with $\mathrm{scal}_{g_{N}}\geq\sigma$, $\sigma\in\mathbb{R}$, and let $\Omega\subseteq N$ with $M=\partial\Omega$. We set
$\textbf{m}_{\sigma}(M)=\Lambda(M,g_{M},\sigma)-\int_{M}H_{g}{\rm dvol}_{g_{M}}.$
This should be compared with Bray’s inner mass [10] which also relies on fill-ins. Theorem 1.2 immediately yields
@@
Let $M$ be spin. Then the quasi-local mass $\textbf{m}_{\sigma}(M)$ is well-defined. Moreover, $\textbf{m}_{\sigma}(M)\geq 0$.
Here $\tfrac{1}{2}\sigma$ corresponds to the cosmological constant. We point out that the ADM mass for asymptotically flat manifolds has been successfully extended to the setting $\sigma<0$ [29], though there is no analogue for $\sigma>0$ [12]. On the other hand $\textbf{m}_{\sigma}$ is well-defined and non-negative for any choice of $\sigma\in\mathbb{R}$.
When $n=3$ and $M\cong S^{2}$, this recovers the notion introduced by Mantoulidis–Miao [22, Theorem 1.5]. As in their setting, we obtain:
@@
Suppose that $n=3$, $M=S^{2}$ and $K(M)>0$ for the Gaussian curvature. Then
@@ Appendix A Collection of auxiliary results
The next result is well-known and we will omit its proofs.

### v2

@@ 8. A new quasi-local mass
Our results have applications in Physics. In 1982, Penrose [Pen82] listed important open problems in General Relativity. Finding a suitable notion of quasi-local mass was listed first.
@@
In order to define the mass of a finite region–known as quasi-local mass–various proposals have been made, including the Hawking mass [Haw68, HI01], the Bartnik mass [Bar89], the Brown–York mass [BY93, ST02], and the Wang–Yau mass [WY09]. Each of these notions has attractive geometric features but also important limitations. A striking common feature is that all of them are particularly well-suited for spherical topology: The Hawking mass often underestimates the true mass and is most effectively used together with inverse mean curvature flow, whose monotonicity holds only for $M\cong S^{2}$. The Brown–York and Wang–Yau masses rely on isometric embeddings into Euclidean or Minkowski space, and classical embedding results such as the Weyl theorem require positive Gauss curvature. The Bartnik mass, perhaps the most mathematically natural proposal, has so far been explicitly computed only for spherical boundaries.
This contrasts with the increasing importance of spacetimes with more complicated topology including Emparan–Reall’s black ring [ER02], the Kottler metric [LN15], and the Horowitz–Myers soliton [BH24].
Motivated by [MM17] we define a new notion of quasi-local mass which has no such topological shortcomings. For this purpose, let $(N,g_{N})$ be a complete oriented manifold with $\mathrm{scal}_{g_{N}}\geq\sigma$, $\sigma\in\mathbb{R}$, and let $M\subset N$ be a mean convex hypersurface and let $g_{M}$ be the metric on $M$ induced by $g_{N}$. We set
$\textbf{m}_{\sigma}(M\subseteq(N,g_{N}))=\Lambda(M,g_{M},\sigma)-\int_{M}H_{g}{\rm dvol}_{g_{M}},$
where we define
[…]
to be the “optimal” constant for which Theorem (B) holds. This should be compared with Bray’s inner mass [Bra01] which also relies on fill-ins. Theorem (B) immediately yields
@@
Suppose that $\operatorname{dim}(N)\leq 7$ and $M$ is spin. Then the quasi-local mass $\textbf{m}_{\sigma}(M\subseteq(N,g_{N}))$ is well-defined and $\textbf{m}_{\sigma}(M\subseteq(N,g_{N}))\geq 0$.
Here $\tfrac{1}{2}\sigma$ corresponds to the cosmological constant. We point out that the ADM mass for asymptotically flat manifolds has been successfully extended to the setting $\sigma<0$ [Wan01], though there is no analogue for $\sigma>0$ [BMN11]. On the other hand $\textbf{m}_{\sigma}$ is well-defined and non-negative for any choice of $\sigma\in\mathbb{R}$.
When $n=2$ and $M\cong S^{2}$, this recovers the notion introduced by Mantoulidis–Miao [MM17, Theorem 1.5]. As in their setting, we obtain:
@@
Suppose that $n=2$, $M=S^{2}$ and $K(M)>0$ for the Gaussian curvature. Then
@@ Appendix A Collection of auxiliary results
The next result is well-known and we will omit its proof.
@@
Appendix B Example of unbounded total mean curvature
Let $n\geq 2$. Then, for every $K>0$, there exists a metric $g_{K}$ on $S^{n}$ satisfying
  • $\mathrm{vol}(S^{n},g_{K})\leq K^{-1}$,
  • $\mathrm{diam}(S^{n},g_{K})\leq K^{-1}$,
  • there is a fill-in of $(S^{n},g_{K})$ with nonnegative scalar curvature, positive mean curvature and total mean curvature larger than $K$.
Hence, $\Lambda(S^{n},g_{K},0,0)\geq K$.
By taking connected sums with $g_{K}$, we can implant these examples into an arbitrary Riemannian manifold. More precisely, let $(M,g_{M})$ be a Riemannian manifold and let $(\Omega,g_{\Omega})$ be a fill in with nonnegative scalar and mean curvature and
$\int_{\partial\Omega}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}=\Lambda.$
Then, for $K>0$, there exists a metric $g_{M,K}$ on $M$ and a fill-in $(\Omega_{\varepsilon},g_{\Omega,\varepsilon})$ of $(M,g_{M,K})$ with nonnegative scalar curvature and positive mean curvature, which satisfies
  1. (i)
    $g_{M}|_{U}=g_{M,K}|_{U}$ for a subset $U\subset M$ with $\mathrm{vol}(U)=\mathrm{vol}(M)-K^{-1}$,
  2. (ii)
    $\mathrm{vol}(M,g_{M,K})\leq\mathrm{vol}(M,g_{M})+K^{-1}$,
  3. (iii)
    $\mathrm{diam}(M,g_{M,K})\leq{\rm diam}(M,g_{M})+K^{-1}$,
  4. (iv)
    $\int_{\partial\Omega}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}=\Lambda+K$.
To construct the metric $g_{K}$ on $S^{n}$, let $\ell\coloneqq(2c(n)K)^{2n-1}$ and $r\coloneqq(2c(n)K)^{-2}$, where $c(n)=\mathrm{vol}(S^{n},g_{\circ})$ is a dimensional constant. By increasing $K$, we may assume that $\ell\in\mathbb{N}$ and $20r\leq K^{-1}$. We embed $\ell+1$ disjoint spheres of radius $r$ into $\mathbb{R}^{n+1}$, such that one of them is centered at the origin and the straight paths connecting a given embedded sphere with this “center sphere” does not intersect any other sphere. We perform connected sums to connect all outer spheres to the center-sphere by straight tubes, and we may assume that any connecting tube has volume smaller than $\tfrac{1}{2\ell}K^{-1}$. Let us denote the resulting Riemannian manifold by $(X,g_{X})$. We observe, that $X\approx S^{n}$. By [LM84, Theorem 3.1], we may assume that $X$ has positive mean curvature with respect to the euclidean metric, since $n\geq 2$. Hence, the interior of $X$ constitutes a flat fill-in $(\widetilde{\Omega},g_{\rm flat})$ with positive mean curvature. A visualisation can be found in Fig. 12.
Figure 12. The metric $g_{K}$ on $S^{n}$.
Let us now estimate the volume and the total mean curvature of this fill-in. For the total mean curvature, we can restrict to the respective outer hemispheres, where the mean curvature is constantly equal to $n/r$. Hence, we obtain:
(B.1) $\begin{split}\int_{\partial\widetilde{\Omega}}H_{g_{\rm flat}}\mathrm{dvol}_{g_{\partial\widetilde{\Omega}}}\geq{}&\ell\int_{S^{n}_{r,+}}H_{g_{\rm flat}}\mathrm{dvol}_{g_{\partial\widetilde{\Omega}}}\geq\frac{1}{2}nc(n)\ell r^{n-1}\\ \geq{}&\frac{1}{2}nc(n)\bigl(2c(n)K\bigr)^{2n-1}\bigl(2c(n)K\bigr)^{-2(n-1)}\geq K.\end{split}$
The volume on the other hand satisfies:
$\mathrm{vol}(X)$ $\leq\ell\left(\mathrm{vol}(S^{n}_{r})+\frac{1}{2\ell}K^{-1}\right)=c(n)\bigl(2c(n)K\bigr)^{2n-1}\bigl(2c(n)K\bigr)^{-2n}+\frac{1}{2}K^{-1}$
$\leq\frac{1}{2}K^{-1}+\frac{1}{2}K^{-1}=K^{-1}$
Note, that while we cannot put the outer spheres onto a circle of arbitrarily small radius, we can still cut away portions of the tubes as non-embedded Riemannian manifolds, resulting in tubes of length less than $r\leq\tfrac{1}{12}K^{-1}$. The resulting manifold and its fill-in might not be isometric to a Riemannian submanifold of $(\mathbb{R}^{n+1},\delta)$, but the mean curvature will still be positive, the scalar curvature will remain flat and estimate (B.1) is still valid, as we do not alter the outer hemispheres by this procedure. The volume may have further decreased and the diameter of the resulting metric $g_{K}$ is smaller than $3\pi\cdot r+4\cdot r\leq 12\cdot r\leq K^{-1}$.